# Fluctuations for non-Hermitian dynamics

### v1

@@ 4 Independence of singular values
$+\frac{\zeta}{2N}\sum_{k}\psi^{\prime\prime}(x_{k})v_{k}^{2}{\rm d}\langle b_{k}\rangle_{s}$ (4.27)
$+\frac{\zeta^{2}}{2N}\sum_{k}\psi^{\prime}(x_{k})^{2}v_{k}^{2}{\rm d}\langle b_{k}\rangle_{s}$ (4.28)
@@
$-\sum_{k}2\zeta\psi^{\prime}(x_{k})\phi_{k}v_{k}\frac{{\rm d}\langle b_{k},\xi_{k}\rangle_{s}}{\sqrt{2N}}+\sum_{k}{\rm d}\left\langle\phi_{k}\xi_{k}-\zeta\psi^{\prime}(x_{k})v_{k}\frac{b_{k}}{\sqrt{2N}}\right\rangle_{s}$ (4.30)
$+\sum_{k}\phi_{k}^{2}{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}.$ (4.31)
@@
For $(j,k)\in I_{\mathcal{S}}$ there are two cases: either $\min\{|j|,|k|\}\leqslant cN,|j-k|\leqslant\ell$ or $\min\{|j|,|k|\}\geqslant cN$. From $\|\psi^{\prime}\|_{\infty}\leqslant 1$ and the assumption $\zeta=N/\ell$, in the first case we have $\zeta|\psi(x_{k})-\psi(x_{j})|\leqslant\zeta|x_{k}-x_{j}|\leqslant\tfrac{|k-j|}{c_{G}\ell}=\operatorname{O}(1)$, where we have used that for $t\leqslant\tau$ from (4.14) we have $|x_{k}(t)-x_{j}(t)|\leqslant|k-j|/(c_{G}N)$ when $|k|,|j|\leqslant c_{G}N$, and we choose $c$ small enough compared to $c_{G}$ (say $c=c_{G}/2$). This implies that $\left|\frac{\phi_{k}}{\phi_{j}}+\frac{\phi_{j}}{\phi_{k}}-2\right|=\operatorname{O}(1)\zeta^{2}|x_{k}-x_{j}|^{2}$. In the second case, $|x_{k}|,|x_{j}|\geqslant c/c_{G}$, and $\zeta|\psi(x_{k})-\psi(x_{j})|=0$. One concludes easily that (4.26) is $\operatorname{O}(1)\zeta^{2}(\ell/N)X_{s}$. The terms (4.27) and (4.28) are of smaller order by $\|\psi^{\prime}\|_{\infty}\leqslant 1$, $\|\psi^{\prime\prime}\|_{\infty}\leqslant\zeta$ and $\tfrac{{\rm d}}{{\rm d}t}\langle b_{k}\rangle\leqslant 1$.
@@
$2\frac{\zeta}{N}\sum_{|j|\leqslant N^{1-2\nu},|k|\geqslant N^{1-\nu}}\frac{\psi^{\prime}(x_{j})v_{j}^{2}}{x_{j}-x_{k}}\leqslant\zeta\sum_{j}v_{j}^{2}=\zeta X_{s}\leqslant\frac{\xi}{N^{2}}=\frac{1}{\ell N}.$
@@
$\sum_{k}\frac{\zeta\phi_{k}v_{k}}{2N}+\frac{\zeta^{2}v_{k}^{2}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}=\sum_{k}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+\frac{\zeta^{2}X_{s}}{2N}.$
@@
$\sum_{k}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$ $=\sum_{|k|\leqslant K^{\prime}}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+\sum_{|k|\geqslant K^{\prime}}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$
$=\sum_{|k|\leqslant K^{\prime}}\left(\frac{\zeta v_{k}}{2N}+\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+e^{-K^{\prime}/\ell}\sum_{k}\left(\frac{\zeta v_{k}}{2N}+\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$
$\lesssim\frac{\zeta K^{\prime}}{N}\sqrt{\sum_{|k|\leqslant K^{\prime}}\frac{v^{2}_{k}}{2K^{\prime}}}+\sum_{|k|\leqslant K^{\prime}}\frac{N^{-\widetilde{\omega}}}{2N}+e^{-K^{\prime}/\ell}\left(\zeta\sqrt{\frac{\sum_{k}v_{k}^{2}}{N}}+\sum_{k}\frac{1}{N}\right)$
$\lesssim\frac{1}{N^{1+\omega_{\ell}/4}}+\frac{1}{N^{1+\omega_{K}-\widetilde{\omega}}}$
where in the third line we used that for $|k|\leqslant K$, ${\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}/{\rm d}s\lesssim N^{-\widetilde{\omega}}/N$; in the last line we used that for $s\leqslant\tau$, $\sum_{k}v_{k}^{2}(s)=X_{s}\leqslant 1/N^{2}$, and $\sqrt{K^{\prime}}/\ell\leqslant N^{-\omega_{\ell}/4}$.
@@
$\frac{{\rm d}}{{\rm d}s}\left\langle\sum_{k}-\xi\psi^{\prime}(x_{k})v_{k}^{2}\frac{b_{k}(s)}{\sqrt{2N}}\right\rangle\lesssim\frac{\xi^{2}}{N}\left(\sum_{k}v_{k}^{2}\right)^{2}=\frac{\xi^{2}X^{2}_{s}}{N}\lesssim\frac{1}{\ell^{2}N^{3+4\omega}},$
@@
$\frac{{\rm d}}{{\rm d}s}\left\langle\sum_{k}v_{k}\phi_{k}\frac{\xi_{k}}{\sqrt{2N}}\right\rangle$ $\leqslant\sum_{k,j\leqslant K}v_{k}v_{j}\phi_{k}\phi_{j}\frac{N^{-\tilde{\omega}}}{2N}+\sum_{k\geqslant K\text{ or }j\geqslant K}v_{k}v_{j}\phi_{k}\phi_{j}\frac{C}{2N}$
$\leqslant\frac{1}{N^{1+\tilde{\omega}-\omega_{K}}}\sum_{k}v_{k}^{2}+e^{-K/\ell}\sum_{k}v_{k}^{2}\lesssim\frac{1}{N^{1+\tilde{\omega}-\omega_{K}}}X_{s}\leqslant\frac{1}{N^{3+\tilde{\omega}+2\omega-\omega_{K}}}.$
@@
$\sup_{t\leqslant\tau}\left|\int_{0}^{s}\sum_{k}-\xi\psi_{k}^{\prime}(u)v_{k}^{2}(u)\frac{{\rm d}b_{k}(u)}{\sqrt{2N}}\right|\leqslant\frac{\sqrt{t}}{\ell N^{3/2}},\quad\sup_{t\leqslant\tau}\left|\int_{0}^{s}\sum_{k}v_{k}\psi_{k}\frac{{\rm d}\xi_{k}}{\sqrt{2N}}\right|\leqslant\frac{\sqrt{t}}{N^{3/2}N^{(\tilde{\omega}-\omega_{K})/2}}.$
@@
$X_{s}$ $\leqslant C\left(\frac{\log Nt}{N\ell}+\frac{\sqrt{t}}{\ell N^{3/2}}+\frac{\sqrt{t}}{N^{3/2}N^{(\tilde{\omega}-\omega_{K})/2}}+\frac{t}{N^{1+\omega_{\ell}/4}}+\frac{t}{N^{1+\omega_{K}-\widetilde{\omega}}}\right)$
$\leqslant\frac{C}{N^{2}}\left(\frac{1}{N^{\omega_{\ell}-\omega_{t}}}+\frac{1}{N^{\omega_{\ell}-\omega_{t}/2}}+\frac{1}{N^{(\tilde{\omega}-\omega_{K}-\omega_{t})/2}}+\frac{1}{N^{\omega_{\ell}/4-\omega_{t}}}+\frac{1}{N^{\omega_{K}-\widetilde{\omega}-\omega_{t}}}\right)\ll\frac{1}{N^{2+2\omega}},$
where we used $\omega\ll\omega_{t}\ll\omega_{\ell}<\omega_{K}\ll\tilde{\omega}$. We conclude that with very high probability $\tau=t$, and $X_{t}\ll N^{-2-2\omega}$. Thanks to the choice of the stopping time (4.32), $|x_{i}(t,\alpha)-x_{j}(t,\alpha)|\leqslant|i-j|\varphi/N$. Thus for any $|i|\leqslant\ell$, $|x_{i}(t,\alpha)|\lesssim\ell/N$, and $\phi_{i}(t)\asymp 1$. We conclude that $N^{-2-2\omega}\geqslant X_{t}\gtrsim w_{i}(t)^{2}$, so that $|w_{i}(t)|\leqslant N^{-1-\omega}$ with very high probability. ∎
@@
$\mu_{t}=\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\delta_{x_{i}(t)},\qquad\quad m_{t}(z)=\frac{1}{2N}\sum_{i}\frac{1}{x_{i}(t)-z}.$
We assume that the initial data $\mu_{0}$ is $(\nu,G)$-regular around $\widetilde{\mu}_{0}$ as in Assumption 4.3.
@@
Let $\widetilde{\mu}_{t}=\widetilde{\mu}_{0}\boxplus\mu_{\rm sc}^{(t)}$ denote the free convolution between $\widetilde{\mu}_{0}$ with the rescaled semicircle distribution. By [biane1997free, Corollary 2], $\widetilde{\mu}_{t}$ has a density $\widetilde{\rho}_{t}:\mathbb{R}\rightarrow\mathbb{R}_{\geqslant 0}$ with respect to Lebesgue measure for $t>0$. Following [biane1997free], the Stieltjes transform $\widetilde{m}_{t}$ of $\widetilde{\mu}_{t}$ is characterized in the following way. For any $t>0$, denote the function $\Phi=\Phi^{t}:\mathbb{H}\rightarrow\mathbb{C}$ and the set $\Lambda_{t}\subseteq\mathbb{H}$ by (see Figure 1)
@@
From [biane1997free, Lemma 4], the function $\Phi$ is a homeomorphism from $\overline{\Lambda}_{t}$ to $\overline{\mathbb{H}}$. Moreover, it is a holomorphic map from $\Lambda_{t}$ to $\mathbb{H}$ and a bijection from $\partial\Lambda_{t}$ to $\mathbb{R}$. For any real number $t\geqslant 0$, $\widetilde{m}_{t}=:\mathbb{H}\rightarrow\mathbb{H}$ is characterized as
@@
Let ${\mathsf{T}}=\varphi^{-2}$ and recall the definition of $\varphi$ in (4.1). Assume that $\mu_{0}$ is $(\nu,G)$ regular. Then, we have the following local law with high probability for any $t\in[0,{\mathsf{T}}]$
$\left|m_{t}(w)-\widetilde{m}_{t}(w)\right|\leqslant\varphi\sqrt{\frac{{\rm Im}[\widetilde{m}_{0}(w)]}{N{\rm Im}[w]}},\quad w\in\{z\in{\mathbb{H}}:|z|\leqslant G/2,{\rm Im}[z]\geqslant\varphi^{4}N^{-1+\nu}\}.$
@@
The first statement in Lemma 4.8 follows from the lower and upper bound of $m_{t}(w)$, which follows from Proposition 4.13. We omit its proof and refer to [erdHos2009local, Theorem 1.1] for the proof.
To prove (4.15), we interpolate (4.35) with the standard Brownian motion, for which we know the optimal rigidity, i.e. with very high probability the $i$-th particle is close to the classical locations with error $\varphi/N$. Thanks to Lemma 4.10, the interpolation induces an error bounded by $\varphi\sqrt{t/N}$. It follows that
$|x_{i}(t)-\gamma_{i}(t)|\lesssim\varphi\left(\frac{1}{N}+\sqrt{\frac{t}{N}}+t\right)\lesssim\varphi\left(\frac{1}{N}+t\right),$
and (4.15) follows, by noticing that the classical locations $\gamma_{i}(t)$ moves with speed $\operatorname{O}(t)$. ∎
@@
${\mathcal{D}}_{t}\mathrel{\mathop{:}}=\left\{z\in{\mathbb{H}}:|z|\leqslant G-Ct,{\rm Im}[z]\,{\rm Im}[\widetilde{m}_{t}(z)]\geqslant\varphi^{4}/N\right\}$ (4.39)
@@
From our assumption for $|u|\leqslant G$, $|\widetilde{m}_{0}(u)|\leqslant C$, and $|z_{t}(u)-u|=t|\widetilde{m}_{0}(u)|$. Thus $z_{t}(u)$ maps $\{u:|u|\leqslant G\}$ surjectively to $\{z:|z|\leqslant G-Ct\}$. We also conclude that for $z=z_{t}(u)$ with $|z|\leqslant G-Ct$, it holds $\widetilde{m}_{t}(z)=\widetilde{m}_{0}(u)$, and (4.40) holds from the definition of ${\mathcal{D}}_{t}$. Moreover, for any $z\in{\mathcal{D}}_{t}$, we have ${\rm Im}[\widetilde{m}_{t}(z)]\geqslant 1/C$. It is easy to see that ${\mathcal{D}}_{t}$ is not empty, and in fact $\{z\in{\mathbb{H}}:|z|\leqslant G/2,{\rm Im}[z]\geqslant C\varphi^{4}N^{-1}\}\subset{\mathcal{D}}_{t}$. The statement (4.40) follows from the definition of ${\mathcal{D}}_{t}$ from(4.39).
@@
=$\mathcal{L}=\left\{E+{\rm i}\eta\in\operatorname{\mathcal{D}}_{0}:E\in{\mathbb{Z}}/N^{6},\eta\in{\mathbb{Z}}/N^{6}\right\}.$ (4.42)
@@
It follows from Lemma 4.14 that if $z_{t}(u)\in\operatorname{\mathcal{D}}_{t}$, then $u\in{\mathcal{D}}_{0}$, and $z_{t}(u)=u-tm_{0}(u)$ with $u\in\operatorname{\mathcal{D}}_{0}$. In particular $|\partial_{u}z_{t}(u)|=|1-tm_{0}^{\prime}(u)|\leqslant 1+C{\mathsf{T}}$. Thus $z_{t}(u)$ is Lipschitz in $u$ with Lipschitz constant bounded by $(1+C{\mathsf{T}})$. Thus for any $w\in\operatorname{\mathcal{D}}_{t}$, there exists some lattice point $u\in\mathcal{L}\cap z_{t}^{-1}(\operatorname{\mathcal{D}}_{t})$, such that $|z_{t}(u)-w|\leqslant(1+C{\mathsf{T}})/N^{6}\leqslant N^{-5},$ provided $N$ is large enough. ∎
@@
${\rm d}m_{s}(z)=m_{s}(z)\partial_{z}m_{s}{\rm d}s+\frac{1}{4N^{2}}\sum_{|k|\leqslant N}\frac{{\rm d}\langle B_{k}\rangle_{s}-{\rm d}s}{(x_{k}(s)-z)^{3}}-\frac{1}{2\sqrt{2}N^{3/2}}\sum_{|k|\leqslant N}\frac{{\rm d}B_{k}(s)}{(z-x_{k}(s))^{2}}.$ (4.43)

### v2

@@ 4 Independence of singular values
$+\frac{\zeta}{2N}\sum_{1\leqslant|k|\leqslant N}\psi^{\prime\prime}(x_{k})v_{k}^{2}{\rm d}\langle b_{k}\rangle_{s}$ (4.30)
$+\frac{\zeta^{2}}{2N}\sum_{1\leqslant|k|\leqslant N}\psi^{\prime}(x_{k})^{2}v_{k}^{2}{\rm d}\langle b_{k}\rangle_{s}$ (4.31)
@@
$-\sum_{1\leqslant|k|\leqslant N}2\zeta\psi^{\prime}(x_{k})\phi_{k}v_{k}\frac{{\rm d}\langle b_{k},\xi_{k}\rangle_{s}}{\sqrt{2N}}+\sum_{1\leqslant|k|\leqslant N}{\rm d}\left\langle\phi_{k}\xi_{k}-\zeta\psi^{\prime}(x_{k})v_{k}\frac{b_{k}}{\sqrt{2N}}\right\rangle_{s}$ (4.33)
$+\sum_{1\leqslant|k|\leqslant N}\phi_{k}^{2}{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}.$ (4.34)
@@
For $(j,k)\in I_{\mathcal{S}}$ there are two cases: either $\min\{|j|,|k|\}\leqslant cN,|j-k|\leqslant\ell$ or $\min\{|j|,|k|\}\geqslant cN$. From $\|\psi^{\prime}\|_{\infty}\leqslant 1$ and the assumption $\zeta=N/\ell$, in the first case we have $\zeta|\psi(x_{k})-\psi(x_{j})|\leqslant\zeta|x_{k}-x_{j}|\leqslant\tfrac{|k-j|}{c_{R}\ell}=\operatorname{O}(1)$, where we have used that for $t\leqslant\tau$ from (4.17) we have $|x_{k}(t)-x_{j}(t)|\leqslant|k-j|/(c_{R}N)$ when $|k|,|j|\leqslant c_{R}N$, and we choose $c$ small enough compared to $c_{R}$ (say $c=c_{R}/2$). This implies that $\left|\frac{\phi_{k}}{\phi_{j}}+\frac{\phi_{j}}{\phi_{k}}-2\right|=\operatorname{O}(1)\zeta^{2}|x_{k}-x_{j}|^{2}$. In the second case, $|x_{k}|,|x_{j}|\geqslant c/c_{R}$, and $\zeta|\psi(x_{k})-\psi(x_{j})|=0$. One concludes easily that (4.29) is $\operatorname{O}(1)\zeta^{2}(\ell/N)X_{s}$. The terms (4.30) and (4.31) are of smaller order by $\|\psi^{\prime}\|_{\infty}\leqslant 1$, $\|\psi^{\prime\prime}\|_{\infty}\leqslant\zeta$ and $\tfrac{{\rm d}}{{\rm d}t}\langle b_{k}\rangle\leqslant 1$.
@@
$2\frac{\zeta}{N}\sum_{|j|\leqslant N^{1-2\nu},|k|\geqslant N^{1-\nu}}\frac{\psi^{\prime}(x_{j})v_{j}^{2}}{x_{j}-x_{k}}\leqslant\zeta\sum_{1\leqslant|j|\leqslant N}v_{j}^{2}=\zeta X_{s}\leqslant\frac{\xi}{N^{2}}=\frac{1}{\ell N}.$
@@
$\sum_{1\leqslant|k|\leqslant N}\frac{\zeta\phi_{k}v_{k}}{2N}+\frac{\zeta^{2}v_{k}^{2}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}=\sum_{1\leqslant|k|\leqslant N}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+\frac{\zeta^{2}X_{s}}{2N}.$
@@
$\sum_{1\leqslant|k|\leqslant N}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$ $=\sum_{|k|\leqslant K^{\prime}}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+\sum_{|k|\geqslant K^{\prime}}\left(\frac{\zeta\phi_{k}v_{k}}{2N}+\phi_{k}^{2}\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$
$=\sum_{|k|\leqslant K^{\prime}}\left(\frac{\zeta v_{k}}{2N}+\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)+e^{-K^{\prime}/\ell}\sum_{1\leqslant|k|\leqslant N}\left(\frac{\zeta v_{k}}{2N}+\frac{{\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}}{{\rm d}s}\right)$
$\lesssim\frac{\zeta K^{\prime}}{N}\sqrt{\sum_{|k|\leqslant K^{\prime}}\frac{v^{2}_{k}}{2K^{\prime}}}+\sum_{|k|\leqslant K^{\prime}}\frac{N^{-\widetilde{\omega}}}{2N}+e^{-K^{\prime}/\ell}\left(\zeta\sqrt{\frac{\sum_{1\leqslant|k|\leqslant N}v_{k}^{2}}{N}}+\sum_{1\leqslant|k|\leqslant N}\frac{1}{N}\right)$
$\lesssim\frac{1}{N^{1+\omega_{\ell}/4}}+\frac{1}{N^{1-\omega_{K}+\widetilde{\omega}}}$,
where in the third line we used that for $|k|\leqslant K$, ${\rm d}\langle\xi_{k},\xi_{k}\rangle_{s}/{\rm d}s\lesssim N^{-\widetilde{\omega}}/N$; in the last line we used that for $s\leqslant\tau$, $\sum_{1\leqslant|k|\leqslant N}v_{k}^{2}(s)=X_{s}\leqslant 1/N^{2}$, and $\sqrt{K^{\prime}}/\ell\leqslant N^{-\omega_{\ell}/4}$.
@@
$\frac{{\rm d}}{{\rm d}s}\left\langle\sum_{1\leqslant|k|\leqslant N}-\xi\psi^{\prime}(x_{k})v_{k}^{2}\frac{b_{k}(s)}{\sqrt{2N}}\right\rangle\lesssim\frac{\xi^{2}}{N}\left(\sum_{1\leqslant|k|\leqslant N}v_{k}^{2}\right)^{2}=\frac{\xi^{2}X^{2}_{s}}{N}\lesssim\frac{1}{\ell^{2}N^{3+4\omega}},$
@@
$\frac{{\rm d}}{{\rm d}s}\left\langle\sum_{1\leqslant|k|\leqslant N}v_{k}\phi_{k}\frac{\xi_{k}}{\sqrt{2N}}\right\rangle$ $\leqslant\sum_{|k|,|j|\leqslant K}v_{k}v_{j}\phi_{k}\phi_{j}\frac{N^{-\widetilde{\omega}}}{2N}+\sum_{|k|\geqslant K\text{ or }|j|\geqslant K}v_{k}v_{j}\phi_{k}\phi_{j}\frac{C}{2N}$
$\leqslant\frac{1}{N^{1+\widetilde{\omega}-\omega_{K}}}\sum_{1\leqslant|k|\leqslant N}v_{k}^{2}+e^{-K/\ell}\sum_{1\leqslant|k|\leqslant N}v_{k}^{2}\lesssim\frac{1}{N^{1+\widetilde{\omega}-\omega_{K}}}X_{s}\leqslant\frac{1}{N^{3+\widetilde{\omega}+2\omega-\omega_{K}}}.$
@@
$\sup_{t\leqslant\tau}\left|\int_{0}^{s}\sum_{1\leqslant|k|\leqslant N}-\xi\psi_{k}^{\prime}(u)v_{k}^{2}(u)\frac{{\rm d}b_{k}(u)}{\sqrt{2N}}\right|\leqslant\frac{\sqrt{t}}{\ell N^{3/2}},\quad\sup_{t\leqslant\tau}\left|\int_{0}^{s}\sum_{1\leqslant|k|\leqslant N}v_{k}\psi_{k}\frac{{\rm d}\xi_{k}}{\sqrt{2N}}\right|\leqslant\frac{\sqrt{t}}{N^{3/2}N^{(\widetilde{\omega}-\omega_{K})/2}}.$
@@
[…]
$\leqslant\frac{C}{N^{2}}\left(\frac{1}{N^{\omega_{\ell}-\omega_{t}}}+\frac{1}{N^{\omega_{\ell}-\omega_{t}/2}}+\frac{1}{N^{(\widetilde{\omega}-\omega_{K}-\omega_{t})/2}}+\frac{1}{N^{\omega_{\ell}/4-\omega_{t}}}+\frac{1}{N^{\widetilde{\omega}-\omega_{K}-\omega_{t}}}\right)\ll\frac{1}{N^{2+2\omega}},$
where we used $\omega\ll\omega_{t}\ll\omega_{\ell}<\omega_{K}\ll\widetilde{\omega}$. We conclude that with very high probability $\tau=t$, and $X_{t}\ll N^{-2-2\omega}$. Thanks to the choice of the stopping time (4.35), $|x_{i}(t,\alpha)-x_{j}(t,\alpha)|\leqslant|i-j|\varphi/N$. Thus for any $|i|\leqslant\ell$, $|x_{i}(t,\alpha)|\lesssim\ell/N$, and $\phi_{i}(t)\asymp 1$. We conclude that $N^{-2-2\omega}\geqslant X_{t}\gtrsim w_{i}(t)^{2}$, so that $|w_{i}(t)|\leqslant N^{-1-\omega}$ with very high probability. ∎
@@
$\mu_{t}=\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\delta_{x_{i}(t)},\qquad\quad m_{t}(z)=\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\frac{1}{x_{i}(t)-z}.$
We assume that the initial data $\mu_{0}$ is $(\nu,R)$-regular around $\widetilde{\mu}_{0}$ as in Assumption 4.3.
@@
Let $\widetilde{\mu}_{t}=\widetilde{\mu}_{0}\boxplus\mu_{\rm sc}^{(t)}$ denote the free convolution between $\widetilde{\mu}_{0}$ with the rescaled semicircle distribution. By [biane1997free, Corollary 2], $\widetilde{\mu}_{t}$ has a density $\widetilde{\rho}_{t}:\mathbb{R}\rightarrow\mathbb{R}_{\geqslant 0}$ with respect to Lebesgue measure for $t>0$. Following [biane1997free], the Stieltjes transform $\widetilde{m}_{t}$ of $\widetilde{\mu}_{t}$ is characterized in the following way. For any $t>0$, denote the function $\Phi=\Phi^{t}:\mathbb{H}\rightarrow\mathbb{C}$ and the set $\Lambda_{t}\subseteq\mathbb{H}$ by
@@
From [biane1997free, Lemma 4], the function $\Phi$ is a homeomorphism from $\overline{\Lambda}_{t}$ to $\overline{\mathbb{H}}$. Moreover, it is a holomorphic map from $\Lambda_{t}$ to $\mathbb{H}$ and a bijection from $\partial\Lambda_{t}$ to $\mathbb{R}$ (see Figure 1). For any real number $t\geqslant 0$, $\widetilde{m}_{t}=:\mathbb{H}\rightarrow\mathbb{H}$ is characterized as
@@
Let ${\mathsf{T}}=\varphi^{-2}$ and recall the definition of $\varphi$ in (4.1). Assume that $\mu_{0}$ is $(\nu,R)$ regular. Then, we have the following local law with high probability for any $t\in[0,{\mathsf{T}}]$
$\left|m_{t}(w)-\widetilde{m}_{t}(w)\right|\leqslant\varphi\sqrt{\frac{{\rm Im}[\widetilde{m}_{0}(w)]}{N{\rm Im}[w]}},\quad w\in\{z\in{\mathbb{H}}:|z|\leqslant R/2,{\rm Im}[z]\geqslant\varphi^{4}N^{-1+\nu}\}.$
@@
The first statement (4.18) in Lemma 4.10 follows from the lower and upper bound of $m_{t}(w)$, which follows from Proposition 4.15. We omit its proof and refer to [erdHos2009local, Theorem 1.1] for the proof.
To prove (4.18), we interpolate (4.38) with the standard Dyson Brownian motion,
${\rm d}\widetilde{x}_{i}(t)=\frac{{\rm d}{\widetilde{b}_{i}(t)}}{\sqrt{2N}}+\frac{1}{2N}\sum_{j\neq i}\frac{1}{\widetilde{x}_{i}(t)-\widetilde{x}_{j}(t)}{\rm d}t,\quad 1\leqslant|i|\leqslant N.$ (4.42)
with initial data $\widetilde{x}_{i}(0)=x_{i}(0)$ for $1\leqslant|i|\leqslant N$, where $\widetilde{b}_{1}(t),\widetilde{b}_{2}(t),\cdots,b_{N}(t)$ are independent Brownian motions, and $\widetilde{b}_{-i}(t)=-b_{i}(t)$ for $1\leqslant i\leqslant N$. The evolution (4.42) is the standard BC-type Dyson Brownian motion (singular-value process of matrix-valued Brownian motions), and can be analyzed by the same way as the standard Dyson Brownian motion. In particular the optimal rigidity holds (see [HuaLan2019, Corollary 3.2] and [chelopatto2019, Theorem 4.5]), i.e. with very high probability the $i$-th particle is close to the classical locations with error $\varphi/N$:
$|\widetilde{x}_{i}(t)-\gamma_{i}(t)|\lesssim\frac{\varphi}{N},\qquad\quad\int_{\gamma_{i}(t)}^{\infty}{\rm d}\mu_{t}(x)=\frac{(N-i+1/2)}{2N}.$ (4.43)
We can then interpolate (4.38) with (4.42) as in (4.16), and by Lemma 4.12, the interpolation induces an error bounded by $\varphi\sqrt{t/N}$. We also notice that the classical locations $\gamma_{i}(t)$ move with speed $\operatorname{O}(t)$. It thus follows that
$|x_{i}(t)-x_{i}(0)|\leqslant|\widetilde{x}_{i}(t)-\gamma_{i}(t)|+|x_{i}(t)-\widetilde{x}_{i}(t)|+|\gamma_{i}(t)-\gamma_{i}(0)|\lesssim\varphi\left(\frac{1}{N}+\sqrt{\frac{t}{N}}+t\right)\lesssim\varphi\left(\frac{1}{N}+t\right),$
and (4.18) follows. ∎
@@
${\mathcal{D}}_{t}\mathrel{\mathop{:}}=\left\{z\in{\mathbb{H}}:|z|\leqslant R-Ct,{\rm Im}[z]\,{\rm Im}[\widetilde{m}_{t}(z)]\geqslant\varphi^{4}/N\right\}$ (4.44)
@@
From our assumption for $|u|\leqslant R$, $|\widetilde{m}_{0}(u)|\leqslant C$, and $|z_{t}(u)-u|=t|\widetilde{m}_{0}(u)|$. Thus $z_{t}(u)$ maps $\{u:|u|\leqslant R\}$ surjectively to $\{z:|z|\leqslant R-Ct\}$. We also conclude that for $z=z_{t}(u)$ with $|z|\leqslant R-Ct$, it holds $\widetilde{m}_{t}(z)=\widetilde{m}_{0}(u)$, and (4.45) holds from the definition of ${\mathcal{D}}_{t}$. Moreover, for any $z\in{\mathcal{D}}_{t}$, we have ${\rm Im}[\widetilde{m}_{t}(z)]\geqslant 1/C$. It is easy to see that ${\mathcal{D}}_{t}$ is not empty, and in fact $\{z\in{\mathbb{H}}:|z|\leqslant R/2,{\rm Im}[z]\geqslant C\varphi^{4}N^{-1}\}\subset{\mathcal{D}}_{t}$. The statement (4.45) follows from the definition of ${\mathcal{D}}_{t}$ from (4.44).
@@
$\mathcal{L}=\left\{E+{\rm i}\eta\in\operatorname{\mathcal{D}}_{0}:E\in{\mathbb{Z}}/N^{6},\eta\in{\mathbb{Z}}/N^{6}\right\}.$ (4.47)
@@
It follows from Lemma 4.16 that if $z_{t}(u)\in\operatorname{\mathcal{D}}_{t}$, then $u\in{\mathcal{D}}_{0}$, and $z_{t}(u)=u-tm_{0}(u)$ with $u\in\operatorname{\mathcal{D}}_{0}$. In particular $|\partial_{u}z_{t}(u)|=|1-t\partial_{u}m_{0}(u)|\leqslant 1+C{\mathsf{T}}$. Thus $z_{t}(u)$ is Lipschitz in $u$ with Lipschitz constant bounded by $(1+C{\mathsf{T}})$. Thus for any $w\in\operatorname{\mathcal{D}}_{t}$, there exists some lattice point $u\in\mathcal{L}\cap z_{t}^{-1}(\operatorname{\mathcal{D}}_{t})$, such that $|z_{t}(u)-w|\leqslant(1+C{\mathsf{T}})/N^{6}\leqslant N^{-5},$ provided $N$ is large enough. ∎
@@
${\rm d}m_{s}(z)=m_{s}(z)\partial_{z}m_{s}{\rm d}s+\frac{1}{4N^{2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}\langle b_{i}\rangle_{s}-{\rm d}s}{(x_{i}(s)-z)^{3}}-\frac{1}{2\sqrt{2}N^{3/2}}\sum_{1\leqslant|i|\leqslant N}\frac{{\rm d}b_{k}(s)}{(z-x_{i}(s))^{2}}.$ (4.48)